# Continuum Limits of Lazy Open Quantum Walks

### v1

@@ 3 Continuum Limit of the Lazy Quantum Walk with Decoherence
The previous section introduced the discrete Lindblad formulation of the lazy quantum walk with decoherence. In this section, that framework is applied to specific choices of decoherence models, and their continuum limits are derived to identify the effective dynamics simulated by each case. It is well established that decoherence acting in the position or coin space leads to qualitatively different macroscopic behaviour [López and Paz, 2003, Love and Boghosian, 2005]. The following subsections derive the continuous spacetime limit of lazy quantum walks with both coin and spatial decoherence. Although both models share the same underlying unitary dynamics, they differ in the structure of the Lindblad term, leading to a different effective PDE in the limit.
@@ 3.2 Unitary Expansion and Effective Hamiltonian
$\mathbf{\hat{U}}(k)=\mathbf{\hat{S}}(\varepsilon,k)\,\mathbf{\mathbf{\hat{C}}}(\varepsilon,\theta)$ (24)
@@
$\mathbf{\mathbf{\hat{C}}}(\varepsilon,\theta)=e^{-i\varepsilon\theta G}\approx\mathbb{I}_{3}-i\varepsilon\theta G-\frac{\varepsilon^{2}\theta^{2}}{2}G^{2}+\mathcal{O}(\varepsilon^{3})$ (25)
The shift operator can then be expanded around $\varepsilon$ as:
$\mathbf{\hat{S}}(\varepsilon,k)=e^{-i\varepsilon kJ_{z}^{(3)}}\approx\mathbb{I}_{3}-i\varepsilon kJ_{z}^{(3)}-\frac{\varepsilon^{2}k^{2}}{2}\big(J_{z}^{(3)}\big)^{2}+\mathcal{O}(\varepsilon^{3})$ (26)
Evaluating the first order expansion of $\mathbf{\hat{U}}(k)=\mathbf{\hat{S}}(\varepsilon,k)\mathbf{\mathbf{\hat{C}}}(\varepsilon,\theta)$ small values of $k$ and $\theta$, the unitary operator can be expanded around the identity to find the Hamiltonian $\hat{\mathbf{H}}_{phys}$:
$\mathbf{\hat{U}}(k)\simeq\mathbb{I}_{3}-i\varepsilon\,\hat{\mathbf{H}}_{\text{phys}}+\mathcal{O}(\varepsilon^{2}),$ (27)
@@
$\hat{\mathbf{H}}_{\text{phys}}(k)=k\,J_{z}^{(3)}+\theta\,G=k\!\left(\tfrac{1}{2}\lambda_{3}+\tfrac{\sqrt{3}}{2}\lambda_{8}\right)+\frac{\theta}{3}\big(\lambda_{1}+\lambda_{4}+\lambda_{6}\big),$ (28)
@@
$\hat{\mathbf{H}}_{\text{phys}}=-i\,J_{z}^{(3)}\,\partial_{x}+\theta\,G,$ (29)
@@
$\partial_{t}\psi(x,t)=-i\hat{\mathbf{H}}_{\text{phys}}\psi(x,t)$ (30)
@@
In terms of field theories, the parameter $\theta$ plays the role of an effective mass. This interpretation follows directly from the structure of the Hamiltonian in (29). The kinetic term $iJ_{z}^{(3)}\partial_{x}$ generates advection of the left and right moving components, while the $\theta\,G$ term induces local mixing in coin space that couples these propagation channels. In two-component quantum walks, such mixing terms are responsible for opening an effective mass gap in the dispersion relation. The same mechanism appears here in its $\mathrm{SU}(3)$ generalisation, as the operator $G$ couples the three internal states symmetrically and acts as the analogue of a mass generating interaction. Consequently, the parameter $\theta$ controls the strength of internal-state coupling and determines the opening of an effective mass gap. Small values of $\theta$ yield nearly massless, strongly dispersive behaviour with sharply separated ballistic branches, while larger $\theta$ enhances mixing and progressively suppresses coherent propagation, leading to a flattening of the dispersion relation. This trend is visible in Fig. 2, where increasing the effective mass parameter $\theta$ suppresses ballistic propagation and reduces the maximum distance reached at fixed $N$, whereas smaller $\theta$ corresponds to lighter, more freely propagating dynamics. The Hamiltonian $\hat{\mathbf{H}}_{\text{phys}}$ therefore represents the minimal $\mathrm{SU}(3)$ embedding of a Dirac-type continuum structure, with $J_{z}^{(3)}$ supplying the kinetic generator and $G$ providing the mass coupling.
This structure defines the universal unitary backbone of the model: all differences between the decoherence channels developed below arise exclusively from how their respective Lindblad projectors act on this shared $\mathrm{SU}(3)$ Hamiltonian, while the underlying effective field theory description of the unitary dynamics remains the same.
@@ 3.3 Including Decoherence
$\boldsymbol{\hat{\rho}}(t+\varepsilon)=\boldsymbol{\hat{\rho}}(t)-i\varepsilon[\mathbf{\hat{H}},\rho(t)]+\varepsilon\bar{\gamma}\sum_{j}\!\left(\mathbb{P}_{j}\mathbf{\hat{U}}\boldsymbol{\hat{\rho}}(t)\mathbf{\hat{U}}^{\dagger}\mathbb{P}_{j}^{\dagger}-\tfrac{1}{2}\big\{\mathbf{\hat{U}}^{\dagger}\mathbb{P}_{j}^{\dagger}\mathbb{P}_{j}\mathbf{\hat{U}},\boldsymbol{\hat{\rho}}(t)\big\}\right)$ (31)
@@ 3.3.1 Coin Decoherence
Substituting these projectors into the discrete Lindblad equation and expanding in $\varepsilon$ to first order, we then pass to the continuum and replace $k\mapsto-\,i\,\partial_{x}$ in $\hat{\mathbf{H}}_{\text{phys}}$ and retaining the leading terms in $\varepsilon$. For clarity, it is convenient to introduce the projector onto the diagonal in the coin basis
@@
$\partial_{t}\boldsymbol{\rho}(x,t)\;=\;\,[\boldsymbol{\rho}(x,t),\partial_{x}J_{z}^{(3)}]\;-\;i\,\theta\,[G,\,\boldsymbol{\rho}(x,t)]\;-\;\bar{\gamma}\,D_{c}[\boldsymbol{\rho}(x,t)],$ (35)
@@
Equation (35) is linear, first order in time, and local in $x$, with the spatial derivative entering only through the commutator with $J_{z}^{(3)}$. It preserves Hermiticity and trace, and for $\bar{\gamma}\!\geq\!0$ generates a CPTP semigroup on the coin space at each spatial point. The unitary part retains the full $\mathcal{SU}(3)$ structure of the lazy walk, while the dephasing produces a uniform decay of the six off diagonal Gell–Mann components $\{\lambda_{1},\lambda_{2},\lambda_{4},\lambda_{5},\lambda_{6},\lambda_{7}\}$ and leaves the diagonal sector $(\lambda_{3},\lambda_{8})$ intact. Consequently, the continuum dynamics conserve the coin populations and the overall streaming structure while progressively reducing interference between the left, rest, and right channels in space and time. The absence of any diagonal contributions in $D_{c}$ indicates that the decoherence channel does not alter the populations of the coin states at leading order, but rather damps their mutual phase relations. Physically, this means that probability weight is conserved within each internal level while the interference terms between them decay. As a result, the coherent oscillations characteristic of the unitary walk are gradually suppressed, and the system tends towards a classical mixture in the coin subspace, providing the microscopic mechanism through which classical behaviour emerges in the macroscopic limit.
@@ 3.4 Spatial Decoherence
For completeness, we now consider the complementary case in which decoherence acts on the spatial subspace rather than on the coin. The discrete evolution is still governed by the Lindblad map in Equation (19) with jump operators of the form (20), but the projectors now resolve the position basis:
@@
$L_{x}\;=\;\mathbb{P}_{x}\,\hat{\mathbf{U}}$. (37)
We assume the same weak coupling continuum scaling as in the coin-decoherence case, given by Equations 22 and 23. We then expand the discrete Lindblad step to first order in $\varepsilon$. The discrete Lindblad map is again expanded to first order in $\varepsilon$ and the continuous spacetime limit is taken. The real space description is obtained by inverting the Fourier representation, and consequently the momentum is again replaced by the differential operator $k\mapsto-\,i\,\partial_{x}$ in $\hat{\mathbf{H}}_{\text{phys}}$. To make the structure of the dissipator explicit, it is convenient to view $\boldsymbol{\rho}$ in the position representation, with matrix elements
@@
$\partial_{t}\boldsymbol{\rho}(x,t)\;=\;\,[\boldsymbol{\rho}(x,t),\partial_{x}J_{z}^{(3)}]\;-\;i\,\theta\,[G,\,\boldsymbol{\rho}(x,t)]\;-\;\bar{\gamma}\,D_{x}[\boldsymbol{\rho}(x,t)],$ (41)
@@
Physically this means that probability at each lattice site, including its internal coin distribution, is conserved at leading order, but coherent superpositions between different positions are progressively suppressed. The walk therefore retains its local streaming structure, yet loses the long-range spatial interference that characterises the purely unitary dynamics. In the long-time limit, the state approaches a classical mixture in the position basis, providing a microscopic mechanism for the emergence of spatially classical transport while still resolving the internal coin degrees of freedom at each site. This makes clear why spatial decoherence is typically undesirable when quantum properties are required: dephasing in the position basis destroys the long-range interference that enables quantum advantage, leaving only the coin subspace quantum but removing the mechanism that produces ballistic transport.
@@ 4 Discussion and Conclusion
The coin dephasing Lindblad term removes off-diagonal elements in the coin subspace while leaving spatial coherences between lattice sites intact. Physically, this results in pure coin dephasing. Local internal superpositions decay, but population transport remains coherent and advective. This behaviour is in good agreement with the earlier numerical results of Love and Boghosian [2005], Alberti et al. [2014], which showed that coin noise damps interference patterns but does not fully classicalise the walk. More importantly, this agreement confirms that the DTQW provides a faithful discretisation of the target continuum quantum field theory. The continuum equation represents the macroscopic model of interest, while the underlying discrete dynamics provide a controlled numerical scheme for its simulation. Establishing this correspondence justifies the use of discrete quantum walks as simulators of continuum quantum transport equations, and motivates the systematic derivation of the continuum limit [di Molfetta and Debbasch, 2012]. Spatial dephasing, by contrast, suppresses coherences between different lattice sites while leaving on-site populations unaffected. As a result, phase correlations decay exponentially in space, progressively eliminating long-range interference effects and driving the dynamics toward an incoherent, effectively classical transport regime. This behaviour aligns with previous numerical observations of spatially dephased quantum walks Tude and de Oliveira [2022]. Together, these results show that coin and spatial dephasing impose qualitatively different transport characteristics. Coin dephasing preserves coherent advection with local phase damping, while spatial dephasing removes all spatial coherence and yields incoherent propagation that quickly reaches a classical, diffusive regime. The continuum framework developed here thus bridges microscopic Lindblad dynamics with the macroscopic behaviour observed in open quantum walk simulations.
@@
Appendix A Explicit Derivation of the Grover Coin Generator
[…]
$|u\rangle=\tfrac{1}{\sqrt{3}}(1,1,1)^{\mathsf{T}},\qquad P_{u}=|u\rangle\langle u|,\qquad P_{\perp}=\mathbb{I}_{3}-P_{u}.$
The projectors satisfy
$P_{u}^{2}=P_{u},\qquad P_{\perp}^{2}=P_{\perp},\qquad P_{u}P_{\perp}=0.$
The explicit matrix form is
$P_{u}=\frac{1}{3}\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix},\qquad P_{\perp}=\begin{pmatrix}\frac{2}{3}&-\frac{1}{3}&-\frac{1}{3}\\ -\frac{1}{3}&\frac{2}{3}&-\frac{1}{3}\\ -\frac{1}{3}&-\frac{1}{3}&\frac{2}{3}\end{pmatrix}$.
The Grover coin has eigenvalue +1+1 on $\mathrm{span}\{|u\rangle\}$ and eigenvalue 1-1 on the orthogonal subspace, so it may be written as
$\mathbf{\hat{C}}_{G}=P_{u}-P_{\perp}.$
To express this as an exponential, we define the generator
$G:=P_{\perp}.$
Since $G$ is a projector,
$G^{2}=G,$
and therefore its exponential has the closed form
$e^{-i\theta G}=\mathbb{I}_{3}+(e^{-i\theta}-1)G=P_{u}+e^{-i\theta}P_{\perp}.$
Setting $\theta=\pi$ reproduces the Grover coin:
$e^{-i\pi G}=P_{u}-P_{\perp}=\mathbf{\hat{C}}_{G}.$
For small $\theta$,
$e^{-i\theta G}=\mathbb{I}_{3}-i\theta G+\mathcal{O}(\theta^{2})=\mathbb{I}_{3}-i\theta\!\left(\mathbb{I}_{3}-|u\rangle\langle u|\right)+\mathcal{O}(\theta^{2}),$
which is the form used in the continuum limit expansion.
Appendix B Gell Mann Matrices

### v2

@@ 3 Continuum Limit of the Lazy Quantum Walk with Decoherence
The previous section introduced the discrete Lindblad formulation of the lazy quantum walk with decoherence. In this section, that framework is applied to specific choices of decoherence models, and their continuum limits are derived to identify the effective dynamics simulated by each case. The following subsections derive the continuous spacetime limit of lazy quantum walks with both coin and spatial decoherence. Although both models share the same underlying unitary dynamics, they differ in the structure of the Lindblad term, leading to a different effective PDE in the limit.
@@ 3.2 Unitary Expansion and Effective Hamiltonian
$\mathbf{\tilde{U}}(k)=\mathbf{\tilde{S}}(\varepsilon,k)\,\mathbf{\mathbf{\hat{C}}}(\varepsilon,\bar{\theta}),$ (25)
@@
$\mathbf{\mathbf{\hat{C}}}(\varepsilon,\bar{\theta})=e^{-i\varepsilon\bar{\theta}G}\approx\mathbb{I}_{3}-i\varepsilon\bar{\theta}G-\frac{\varepsilon^{2}\bar{\theta}^{2}}{2}G^{2}+\mathcal{O}(\varepsilon^{3})$ (26)
The shift operator, given by equation (16), can then be expanded around $\varepsilon$ as:
$\mathbf{\tilde{S}}(\varepsilon,k)=e^{-i\varepsilon kJ_{z}^{(3)}}\approx\mathbb{I}_{3}-i\varepsilon kJ_{z}^{(3)}-\frac{\varepsilon^{2}k^{2}}{2}\big(J_{z}^{(3)}\big)^{2}+\mathcal{O}(\varepsilon^{3})$ (27)
Evaluating the first order expansion of $\mathbf{\tilde{U}}(k)=\mathbf{\tilde{S}}(\varepsilon,k)\mathbf{\mathbf{\hat{C}}}(\varepsilon,\bar{\theta})$ small values of $k$ and $\bar{\theta}$, the unitary operator can be expanded around the identity to find the Hamiltonian $\hat{\mathbf{H}}_{phys}$:
$\mathbf{\tilde{U}}(k)\simeq\mathbb{I}_{3}-i\varepsilon\,\hat{\mathbf{H}}_{\text{phys}}+\mathcal{O}(\varepsilon^{2}),$ (28)
@@
$\hat{\mathbf{H}}_{\text{phys}}(k)=k\,J_{z}^{(3)}+\bar{\theta}\,G=k\!\left(\tfrac{1}{2}\lambda_{3}+\tfrac{\sqrt{3}}{2}\lambda_{8}\right)+\frac{\bar{\theta}}{3}\big(\lambda_{1}+\lambda_{4}+\lambda_{6}\big),$ (29)
@@
$\hat{\mathbf{H}}_{\text{phys}}=-i\,J_{z}^{(3)}\,\partial_{x}+\bar{\theta}\,G,$ (30)
@@
$\partial_{t}\psi(x,t)=-i\hat{\mathbf{H}}_{\text{phys}}\psi(x,t),$ (31)
@@
In terms of field theories, the parameter $\theta$ plays the role of an effective mass. This interpretation follows directly from the structure of the Hamiltonian in (30). The kinetic term $iJ_{z}^{(3)}\partial_{x}$ generates advection of the left and right moving components, while the $\bar{\theta},G$ term induces local mixing in coin space that couples these propagation channels. In two-component quantum walks, such mixing terms are responsible for opening an effective mass gap in the dispersion relation. The same mechanism appears here in its $\mathcal{SU}(3)$ generalisation, as the operator $G$ couples the three internal states symmetrically and acts as the analogue of a mass generating interaction. Consequently, the parameter $\bar{\theta}$ controls the strength of internal-state coupling and determines the opening of an effective mass gap. Small values of $\bar{\theta}$ yield nearly massless, strongly dispersive behaviour with sharply separated ballistic branches, while larger $\theta$ enhances mixing and progressively suppresses coherent propagation, leading to a flattening of the dispersion relation. This trend is visible from the $\gamma=0$ lines in Fig. 2, where increasing the effective mass parameter $\bar{\theta}$ suppresses ballistic propagation and reduces the maximum distance reached at fixed $N$, whereas smaller $\bar{\theta}$ corresponds to lighter, more freely propagating dynamics. The Hamiltonian $\hat{\mathbf{H}}_{\text{phys}}$ therefore represents the minimal $\mathcal{SU}(3)$ embedding of a Dirac-type continuum structure, with $J_{z}^{(3)}$ supplying the kinetic generator and $G$ providing the mass coupling.
This structure defines the universal unitary backbone of the model: all differences between the decoherence channels developed below arise exclusively from how their respective Lindblad projectors act on this shared $\mathcal{SU}(3)$ Hamiltonian, while the underlying effective field theory description of the unitary dynamics remains the same.
@@ 3.3 Including Decoherence
$\boldsymbol{\hat{\rho}}(t+\varepsilon)=\boldsymbol{\hat{\rho}}(t)-i\varepsilon[\mathbf{\hat{H}},\rho(t)]+\varepsilon\bar{\gamma}\sum_{j}\!\left(\mathbb{P}_{j}\mathbf{\tilde{U}}\boldsymbol{\hat{\rho}}(t)\mathbf{\tilde{U}}^{\dagger}\mathbb{P}_{j}^{\dagger}-\tfrac{1}{2}\big\{\mathbf{\tilde{U}}^{\dagger}\mathbb{P}_{j}^{\dagger}\mathbb{P}_{j}\mathbf{\tilde{U}},\boldsymbol{\hat{\rho}}(t)\big\}\right)$ (32)
@@ 3.3.1 Coin Decoherence
These projectors are substituted into the discrete Lindblad equation and we expand in $\varepsilon$ to first order. We take the continuum limit and and retain the leading terms in $\varepsilon$. The real space description is obtained by inverting the Fourier representation, and consequently the momentum is replaced by the differential operator $k\mapsto-\,i\,\partial_{x}$ in $\hat{\mathbf{H}}_{\text{phys}}$. For clarity, it is convenient to introduce the projector onto the diagonal in the coin basis
@@
$\partial_{t}\boldsymbol{\rho}(x,t)\;=\;\,[\boldsymbol{\rho}(x,t),\partial_{x}J_{z}^{(3)}]\;-\;i\,\bar{\theta}\,[G,\,\boldsymbol{\rho}(x,t)]\;-\;\bar{\gamma}\,D_{c}[\boldsymbol{\rho}(x,t)],$ (36)
@@
Equation (36) is linear, first order in time, and local in $x$, with the spatial derivative entering only through the commutator with $J_{z}^{(3)}$. It preserves Hermiticity and trace, and for $\bar{\gamma}\!\geq\!0$ generates a CPTP semigroup on the coin space at each spatial point. The unitary part retains the full $\mathcal{SU}(3)$ structure of the lazy walk, while the dephasing produces a uniform decay of the six off diagonal Gell–Mann components $\{\lambda_{1},\lambda_{2},\lambda_{4},\lambda_{5},\lambda_{6},\lambda_{7}\}$ and leaves the diagonal sector $(\lambda_{3},\lambda_{8})$ intact. Consequently, the continuum dynamics preserve the underlying streaming structure while progressively reducing the interference between the left, rest, and right channels in space and time. The absence of diagonal contributions in $D_{c}$ indicates that the decoherence channel does not directly modify the populations of the coin states at leading order, but instead damps their mutual phase coherences. While the unitary evolution may induce oscillations of the coin populations depending on the initial state, the action of the decoherence term selectively suppresses the off-diagonal components of the coin density matrix. Physically, this implies that probability weight is redistributed between internal levels solely through coherent dynamics, whereas interference between coin states is progressively lost. As a result, the coherent oscillations characteristic of the unitary walk are gradually attenuated, and the system evolves towards an effectively classical mixture in the coin subspace, providing a microscopic mechanism for the emergence of classical behaviour in the macroscopic limit.
Figure 3: Discrete time quantum walk versus derived continuum master equation for coin dephasing. Final position probability $P(x,t)$ after $N=100$ steps of the lazy three state quantum walk with $\theta=\pi/2$, shown for several decoherence rates $\gamma$. Solid lines show the solution of the corresponding derived continuum Lindblad equation. The initial state is a spatially delocalised Gaussian wavepacket of width $\sigma_{0}=5$ with a Fourier-symmetric coin superposition $(1,\omega,\omega^{2})/\sqrt{3}$, where $\omega=e^{2\pi i/3}$.
Figure 3 shows the position probability distributions obtained from DTQW and from the corresponding derived continuum equation for several values of the coin dephasing parameter $\gamma$. The figure demonstrates very good agreement between the DTQW and the continuum Lindblad dynamics throughout the range of weak to moderate dephasing considered here ($0\leq\gamma\leq 0.4$), which is the weak decoherence regime in which the continuum master equation is derived. Within this range, the continuum solution accurately reproduces both the overall shape and the width of the position distribution produced by the discrete walk. In particular, the crossover from ballistic, multi peaked propagation at weak dephasing to a smoother, more centrally concentrated profile as the dephasing strength increases is captured faithfully. This agreement provides strong evidence that the DTQW with coin dephasing constitutes a consistent and controlled discretisation of the target continuum open quantum dynamics in its domain of validity.
@@ 3.4 Spatial Decoherence
We now consider the complementary case in which decoherence acts on the spatial subspace rather than on the coin. The discrete evolution is still governed by the Lindblad map in equation (20) with jump operators of the form (21), but the projectors now resolve the position basis:
@@
$L_{x}\;=\;\mathbb{P}_{x}\,\tilde{\mathbf{U}}$. (38)
We assume the same weak coupling continuum scaling as in the coin-decoherence case, given by equations (23) and (24). We then expand the discrete Lindblad step to first order in $\varepsilon$. The discrete Lindblad map is again expanded to first order in $\varepsilon$ and the continuous spacetime limit is taken. Again, the real space description is obtained by inverting the Fourier representation, and consequently the momentum is replaced by the differential operator $k\mapsto-\,i\,\partial_{x}$ in $\hat{\mathbf{H}}_{\text{phys}}$. To make the structure of the dissipator explicit, it is convenient to view $\boldsymbol{\rho}$ in the position representation, with matrix elements
@@
$\partial_{t}\boldsymbol{\rho}(x,t)\;=\;\,[\boldsymbol{\rho}(x,t),\partial_{x}J_{z}^{(3)}]\;-\;i\,\bar{\theta}\,[G,\,\boldsymbol{\rho}(x,t)]\;-\;\bar{\gamma}\,D_{x}[\boldsymbol{\rho}(x,t)],$ (42)
@@
Physically, this implies that the decoherence channel acts locally in position space and does not directly transfer probability between lattice sites or modify the internal coin populations at a given site, but instead suppresses coherent superpositions between different positions. As a consequence, although the walk retains its local streaming dynamics, the loss of long-range spatial coherence alters the interference mechanisms responsible for ballistic transport. The resulting evolution therefore differs qualitatively from the fully unitary walk, leading to a redistribution of probability in space. In the long-time limit, the state approaches a classical mixture in the position basis, providing a microscopic mechanism for the emergence of spatially classical transport while still resolving the internal coin degrees of freedom at each site. This makes clear why spatial decoherence is typically undesirable when quantum properties are required: dephasing in the position basis destroys the long-range interference that enables quantum advantage, leaving only the coin subspace quantum but removing the mechanism that produces ballistic transport.
Figure 4: DTQWs derived continuum master equation for spatial dephasing. Final position probability $P(x,t)$ after $N=100$ steps of the lazy three state quantum walk with $\theta=\pi/2$, shown for several decoherence rates $\gamma$. Solid lines show the solution of the corresponding derived continuum Lindblad equation with spatial dephasing. The initial state is a spatially delocalised Gaussian wavepacket of width $\sigma_{0}=5$ with a Fourier-symmetric coin superposition $(1,\omega,\omega^{2})/\sqrt{3}$, where $\omega=e^{2\pi i/3}$.
Figure 4 shows the final position probability distribution obtained from the DTQW with spatial dephasing (markers) compared to the corresponding continuum Lindblad equation (solid lines), for several decoherence strengths $0\leq\gamma\leq 0.4$. Again, we restrict the range of gamma between these values as the continuum equation is derived under the weak decoherence assumption. As the dephasing rate increases, the dynamics smoothly interpolate from a ballistic, interference dominated regime with pronounced side peaks to a broader, more diffusive profile concentrated around the origin. The continuum model accurately reproduces this overall transition, including the suppression of the ballistic fronts and the emergence of a central peak at larger $\gamma$ values. Small quantitative differences are visible at low decoherence, particularly in the outer regions of the distribution, where lattice scale interference effects remain more pronounced in the discrete walk than in the continuum description. This slightly reduced agreement compared to the coin-dephasing case is expected, since spatial dephasing acts directly on nonlocal coherences between lattice sites, which are more sensitive to coarse graining in the continuum limit. Overall, the plot confirms that the continuum equation captures the dominant transport behaviour induced by spatial decoherence, while smoothing out fine interference features of the underlying discrete dynamics.
@@ 4 Discussion and Conclusion
The coin dephasing Lindblad term removes off diagonal elements in the coin subspace while leaving spatial coherences between lattice sites intact. Physically, this results in pure coin dephasing. Local internal superpositions decay, but population transport remains coherent until diffusion dominates at high $\gamma$ values. This behaviour is in good qualitative agreement with early numerical studies of decoherence in the two state DTQW, which showed that coin noise suppresses interference without immediately destroying ballistic transport [López and Paz, 2003, Brun et al., 2003, Love and Boghosian, 2005, Alberti et al., 2014]. By contrast, spatial decoherence mechanisms were found to rapidly suppress long-range coherence and induce diffusive behaviour [Romanelli et al., 2005], a qualitative distinction that is also observed here. Crucially, the numerical plots demonstrate good agreement between the DTQW and the derived continuum dynamics across the explored parameter regimes. This agreement confirms that the DTQW provides a faithful discretisation of the target continuum quantum field theory. The continuum equation represents the macroscopic model of interest, while the underlying discrete dynamics form a controlled numerical scheme for its simulation. Establishing this correspondence justifies the use of discrete quantum walks as simulators of continuum quantum transport equations and motivates the systematic derivation of the continuum limit [di Molfetta and Debbasch, 2012]. Spatial dephasing, by contrast, suppresses coherences between different lattice sites while leaving on-site populations unaffected. As a result, phase correlations decay exponentially in space, progressively eliminating long-range interference effects and driving the dynamics toward an incoherent, effectively classical transport regime. This behaviour aligns with previous numerical observations of spatially dephased quantum walks Tude and de Oliveira [2022]. Together, these results show that coin and spatial dephasing impose qualitatively different transport characteristics. Coin dephasing preserves coherent advection with local phase damping, while spatial dephasing removes all spatial coherence and yields incoherent propagation that quickly reaches a classical, diffusive regime. The continuum framework developed here thus bridges microscopic Lindblad dynamics with the macroscopic behaviour observed in open quantum walk simulations.
@@
Appendix A Gell Mann Matrices